OpenStudy (anonymous):

when using induction for example if i had to prove say 2 +5 + ... + 3n -1 = n/3 ( n+2).. if it were true would i add n +1 to both sides or n+1 to every n i see?

OpenStudy (anonymous):

n+1 to every n you see

OpenStudy (asnaseer):

The general approach is: 1. Show its true for some base case (in this case you would show its true for n=1) 2. Assume its true for n=k 3. Then prove it must also be true for n=k+1

OpenStudy (anonymous):

ok thank you i have a similar problem just wanted to know how to start it after i prove P1 is true.... just so i understand 3n -1 + n+1 = ((n+1)/3)((n+1)+2)

OpenStudy (asnaseer):

not quite - let me show you step-by-step...

OpenStudy (asnaseer):

you are told to prove that:\[2 +5 + ... + (3n -1) = n( n+2)/3\] 1, Show this is true for n=1 For n=1 we know the last term in the sum is (3*1-1)=(3-1)=2 So the left-hand-side of your identity becomes 2 The right-hand-side for n=1 becomes 1(1+2)/3=3/3=1 In this case the left-hand-side and right-hand-sides do not match therefore this identity is false.

OpenStudy (asnaseer):

I think your identity should be:\[2 +5 + ... + (3n -1) = n( 3n+1)/2\]

OpenStudy (asnaseer):

then, for n=1, the right-hand-side evaluates to 1(3*1+1)/2=4/2=2 which matches the left-hand-side. we can now move onto to step 2...

OpenStudy (asnaseer):

2. Assume its true for n=k This gives us:\[2+5+...+(3k-1)=k(3k+1)/2\] Now we move to step 3...

OpenStudy (anonymous):

yeah i understand how to prove whether it's true or not...but how do start to prove Pn+1...... so 2 + 5+ ... + (3n-1) +n+1 = n(3(n+1)+1)/2 is this what i need to show when proving it on the right side

OpenStudy (asnaseer):

3. Then prove it must also be true for n=k+1 This gives us (on the left-hand-side):\[2+5+...+(3(k+1)-1)\]

OpenStudy (asnaseer):

we now need to manipulate this expression so that we end up with something that looks like our step 2...

OpenStudy (anonymous):

couldn''t you just right it 2+5+...(3k-1) +(k+1)

OpenStudy (asnaseer):

\[2+5+...+(3(k+1)-1)=2+5+...+(3k-1)+(3(k+1)-1)\]what I did here was include the term before the last one

OpenStudy (asnaseer):

now you should spot that all terms up to \((3k-1)\) match our identity in step 2 (which we assumed was true)

OpenStudy (asnaseer):

therefore:\[2+5+...+(3(k+1)-1)=2+5+...+(3k-1)+(3(k+1)-1)\]\[=k(3k+1)/2+(3(k+1)-1)=k(3k+1)/2+(3k+3-1)\]\[=k(3k+1)/2+3k+2\]

OpenStudy (dan815):

why induction for something so simple.. but okay

OpenStudy (asnaseer):

your question about "couldn''t you just right it 2+5+...(3k-1) +(k+1)" is not correct. the identity states:\[2+5+...+(3n-1)=n(3n+1)/2\]this hold for "n" terms. so, if we have "k" terms, we get:\[2+5+...+(3k-1)=n(3k+1)/2\]and if we have "k+1" terms we get:\[2+5+...+(3(k+1)-1)=(k+1)(3(k+1)+1)/2\]

OpenStudy (asnaseer):

@dan815 - I believe the asker wants to learn about induction

OpenStudy (asnaseer):

sorry I forgot to replace one "n" with "k" above ^^^

OpenStudy (asnaseer):

@fiberdust - does this make sense now?

OpenStudy (asnaseer):

basically you have to replace EVERY "n" with "k+1" to get the correct identity for the k+1'th case

OpenStudy (anonymous):

yeah so add k+1 where u see an n every time?

OpenStudy (asnaseer):

not "add" but replace every "n" with "k+1"

OpenStudy (anonymous):

ok thank you all for your help and thoughts for understanding the process better with this type of problem

OpenStudy (asnaseer):

To take a very simple example, lets say you wanted to add all integers up to n, you would have:\[1+2+...+n\]so the sum of all integers up to k would be:\[1+2+...+k\]and the sum of all integers up to k+1 would be:\[1+2+...+(k+1)=1+2+...+k+(k+1)\]

OpenStudy (asnaseer):

yw

OpenStudy (dan815):

is there something wrong,. with the starting equation

OpenStudy (asnaseer):

yes - the identity was not stated correctly

OpenStudy (dan815):

oh okay wow ya

OpenStudy (dan815):

wasttinng my time!!

OpenStudy (dan815):